OpenStudy (anonymous):

8x+2y=-2,y=-5x+1 substitution

OpenStudy (anonymous):

8x+2(-5x+1) = -2 8x-10x+2 = -2 -2x = -4 x = 2

OpenStudy (anonymous):

thankssssssssssss sooooo myuchhhhhh

OpenStudy (anonymous):

you're welcome again lol